The straight-line graph by the equation y=2x-4 will cut the y=axis at the point.


1
Expert's answer
2021-05-11T07:58:55-0400

The equation y=2x-4 will cut the y=axis at y = 2×0 - 4 = -4, so the point is (0; -4).